# A gluing construction of $D_{k}$ ALF gravitational instantons and existence of non-holomorphic minimal spheres

### v1

@@ 2.3. Gluing construction
$\{\frac{\epsilon}{\delta}<|x|<\delta\}\subset M_{\epsilon}$
we use the identification map $x^{\prime}=\kappa(x)=x/\epsilon,\kappa^{*}(\alpha_{0})=\alpha^{\prime}$ to identify $g_{\epsilon}$ and $g_{\widehat{AH},\epsilon}$, and note that under this map $h_{\epsilon}$ and $h_{\epsilon^{\prime}}$ are identified up to $O(\epsilon^{3})$.
However in order to apply the analysis, we need to modify the space. The total space is glued from two parts: $\mathcal{W}_{0}$ corresponding to the Atiyah-Hitchin metrics in the center, and $\mathcal{W}_{1}$ corresponding to the multi-Taub-NUT singularities, where $\mathcal{W}_{0}$ and $\mathcal{W}_{1}$ are essentially $M_{\epsilon}$ and $\widehat{AH}_{\epsilon}$ with the additional steps of compactification at infinity and singularity resolution at $\epsilon=0$. The construction of $\mathcal{W}_{1}$ and $\mathcal{W}_{0}$ are almost the same as in [SS21], while some modification is needed.
The construction is first done on the double cover $\widetilde{\mathcal{W}}_{i},\ i=0,1$, and the final space $\mathcal{W}$ is the quotient by the involution $\iota$. First we construct $\widetilde{\mathcal{W}}_{0}$, which is defined as
@@
The construction of $\widetilde{\mathcal{W}}_{1}$ is the same as in [SS21, Section 4.1], obtained by pulling back the product of interval $[0,\delta^{2})_{\epsilon}\times M_{\epsilon}$ to the resolved base $\widetilde{B}$. We briefly explain the construction of $\widetilde{B}$ here. Define the blow up
(11) $\widetilde{B}=[\bar{\mathbb{R}}^{3}_{x}\times[0,\delta^{2})_{\epsilon};\{(0,0),\pm p_{i}\times\{0\},i=1,\dots,k\}]\cap\{|x|>\epsilon/\delta\}$
which introduced $12k+$ 1 front faces. We then pull back the circle fibration $\phi:M_{\epsilon}\rightarrow\bar{\mathbb{R}}^{3}\setminus\{|x|\leq\epsilon/\delta\}$ to $\widetilde{B}$, such that when restricting to each $\epsilon>0$ it is the same as the original fibration $\phi:M_{\epsilon}\rightarrow\bar{\mathbb{R}}^{3}\setminus\{\frac{\epsilon}{\delta}<|x|\leq\epsilon/\delta\}$. On the other hand, when restricting to $\epsilon=0$ it lifts to a circle bundle over the blown up space [¯$[\bar{\mathbb{R}}^{3};\{0,\pm p_{i}\}]$.
We construct the space gluing $\widetilde{\mathcal{W}}_{1}$ into $\widetilde{\mathcal{W}}_{0}$, identifying the torus neighborhood using the lift of the map $\kappa$ described earlier, i.e. we identify the neighborhood
$\{(z^{\prime},\epsilon^{\prime}):0\leq\epsilon^{\prime}\leq\epsilon_{0},\delta^{-1}<|\psi(z^{\prime})|<\delta\epsilon^{-1}\}\subset\widetilde{\mathcal{W}}_{0}$
@@
(12) $\widetilde{\mathcal{W}}:=\widetilde{\mathcal{W}}_{1}\coprod_{\kappa}\widetilde{\mathcal{W}}_{2},\ \mathcal{W}:=\widetilde{\mathcal{W}}/\iota.$
Here $\mathcal{W}$ is a manifold with corners, where the boundary faces include: $I_{\infty}$ which corresponds to the fibration over the boundary of the compactified $\mathbb{R}^{3}$ of $\mathcal{W}_{1}$; $X_{0}$ that is the lift of the fibration to the boundary face created by blowing up $(0,0)$; $X_{i},i=1,\dots k$, that are the lift of the fibration to the boundary faces created by blowing up {±$\{\pm p_{i},0\}$; $X_{ad}$ the remaining part of $\epsilon=0$. Here is a picture of the total space:
Figure 1. An illustration of the resolved total space $\mathcal{W}$ obtained from gluing $\mathcal{W}_{0}$ and $\mathcal{W}_{1}$, with boundary faces and associated boundary defining functions on $\pi^{-1}(0)$
There is a projection map $\pi:\mathcal{W}\rightarrow[0,\epsilon_{0})_{\epsilon}$, such that for each nonzero $\epsilon$ the fiber is an approximate ALF metric, while for $\epsilon=0$ the fiber is a manifold with corners. We also note that $\epsilon$ is the total boundary defining function, i.e. if we define $\sigma_{\nu}$ to be the boundary defining function of $X_{\nu},\nu=0,1,\dots,k$, and $\rho$ to be the boundary defining function of $X_{ad}$, then the product of all boundary defining functions on the front face is
@@ 2.4. Solving for the hyperKähler triples
By the gluing construction discussed above, we arrived at a family of approximate metric, each living on the fiber $\pi^{-1}(\epsilon)$ of $\mathcal{W}$. We will follow the set up in  [SS21] to perturb these metrics and solve for the exact hyperKähler triples on each fiber. Recall that every hyperKähler metric on a 4-manifold gives a hyperKähler triple $\mathbf{\omega}=(\mathbf{\omega}_{1},\mathbf{\omega}_{2},\mathbf{\omega}_{3})$ such that each $\omega_{i}$ is a symplectic form and they satisfy the following equation [Don06]
@@
For a adiabatic Gibbons-hawking ansatz in (1), the triple is given by the following $\mathbf{\omega}_{\epsilon}=(\mathbf{\omega}_{1},\mathbf{\omega}_{2},\mathbf{\omega}_{3})$ [SS21, (2.21)]
@@
which can be written as $\Omega_{\epsilon}+db$ where $b\in O(|\epsilon|^{3})$ where $\Omega_{\epsilon}$ is given by replacing $h_{\epsilon}$ by its leading term $1-\frac{\epsilon}{|x|}+\sum_{v=1}^{k}\frac{\epsilon}{|p_{v}|}$. Similarly for the branched cover $\widehat{AH}_{\epsilon^{\prime}}$ the triple is given by $\mathbf{\omega}_{AH,\epsilon^{\prime}}=\Omega_{\epsilon^{\prime}}+da$ where $a=O(e^{-|x^{\prime}|})$ and $\Omega_{\epsilon^{\prime}}$ is given by
@@
Fix $\{0,p_{1},\dots,p_{k}\}$ a set of distinct $k+1$ points. Then there exists $\epsilon_{0}>0$, such that on $\pi^{-1}([0,\epsilon_{0}))\subset\mathcal{W}$, there exists smooth triple $\mathbf{\omega}_{Se}=(\mathbf{\omega}_{1},\mathbf{\omega}_{2},\mathbf{\omega}_{3})$ with $\mathbf{\omega}_{i}\in\Omega_{\phi}^{2}(\mathcal{W}/I)$ satisfying $Q(\mathbf{\omega})=0$. The corresponding fiber metric $g_{Se,\epsilon}$ satisfies
@@
    $g_{Se,0}|_{X_{0}}$ restricts to be the Atiyah–Hitchin metric, and $g_{Se,0}|_{X_{i}}$ restricts to be the Taub–NUT metric for $i=1,\dots,k$.
@@
where the two hyperKähler triples are defined in (15) and (16). Note that the harmonic function in $\mathbf{\omega}_{AH,\epsilon^{\prime}}$ and $\mathbf{\omega}_{\epsilon}$ agree up to a closed form of order $\epsilon^{3}$, therefore the initial approximation solves the equation up to the following error
@@
The second step is to solve the equation iteratively to $\epsilon^{\infty}$ error. This step follows from the discussion in [SS21, Section 5], where they solve the error iteratively near the corner of $X_{\nu}\cap X_{ad}$ on the two faces. Recall that the linearized operator of $Q$ is given by 12 copies of scalar Laplacians [SS21]. When restricting to the solvability near $X_{\nu}$, note that the replacement $AH$ is still a strongly ALF metric hence the analysis carries through. The solvability near $X_{ad}$ is the same as before. As a result we can iteratively construct a formal power series solution $\zeta=\mathbf{\omega}_{\chi}+dc$ such that
@@
The third step is to show there is a bounded inverse of the linearized operator on $\mathcal{W}$. This step again relies on that one can patch inverses on $X_{\nu}$ and $X_{ad}$. Since the only difference in the structure of $X_{\nu}$ from [SS21] is that $X_{0}$ is changed from $\overline{M_{0}^{2}}$ by $\overline{AH}$, the inverse is constructed in the same way as in [SS21, Theorem C.3] for a strongly ALF space. On the other hand, on $X_{ad}$ the inverse is constructed explicitly using Fourier series and Green’s function where the only difference is the coefficient in $h_{\epsilon}$ which does not change the proof. As a result $\Delta_{\zeta}$ is an invertible operator on the domain constructed in [SS21, Theorem 6.6].
@@
As a corollary, such a family of metrics degenerates into the front face, and (under correct scaling and choice of base points) give Gromov-Hausdorff convergence to $AH$ and $TN$ respectively.
@@ 3. Non-holomorphic minimal spheres
In [MW93] Micallef and Wolfson showed that in the Atiyah–Hitchin space there is a special sphere that is stably minimal and not holomorphic with respect to any complex structures in the hyperähler rotation. Using the gluing result above, we extend this result to $D_{k+1}$ spaces constructed above. As a perturbation result, the existence can be proved for $D_{k+1}$ spaces where the Taub-NUT singularities are sufficiently far from the origin.
@@
There exists $\epsilon_{0}$, such that for every $\epsilon\in(0,\epsilon_{0})$ there exists a minimal sphere in $g_{Se,\epsilon}$ that is non-holomorphic with respect to any of the compatible complex structures.
@@
We start with the approximate metric $g_{0,\epsilon}$ constructed via patching the $g_{\widehat{AH},\epsilon}$ and $g_{\epsilon}$. Since the minimal sphere in $g_{\widehat{AH}}$ is compact, so by shrinking $\epsilon_{0}$ if necessary, there exists $\Sigma_{\epsilon}\subset(AH|_{|x^{\prime}|<\delta/\epsilon},g_{\widehat{AH},\epsilon})\subset(\pi^{-1}(\epsilon),g_{0,\epsilon})$ for every $\epsilon\in(0,\epsilon_{0})$. Note that this minimal sphere $\Sigma$ satisfies $[\Sigma]\cdot[\Sigma]=-4$ which is not $22*\text{genus}(\Sigma)-$ 2, hence it cannot be holomorphic in a hyperKähler manifold via adjunct formula. Now we apply White’s implicit function theorem [Whi91] (which does not require the ambient manifold to be compact) to the actual hyperKähler metric $g_{SE,\epsilon}$ perturbed from $g_{0,\epsilon}$, we get a stably minimal non-holomorphic sphere in the $D_{k+1}$ space. ∎
@@
There exists a neighborhood of $K3$ such that every $K3$ has a minimal sphere that is non-holomorphic with respect to any of the complex form.
@@
The proof follows exactly the same as in [Fos19, Theorem 7.1] by considering any $D_{k+1}$ metric constructed above as a building block in the collapsing sequence. Since the approximate metric $g_{\epsilon}$ constructed using $D_{k}$ bubbles contains the special sphere, by the same perturbation argument in the previous theorem we obtain the existence of non-holomorphic minimal sphere. ∎

### v2

@@ 2.3. Gluing construction
$\{\frac{\epsilon}{\delta}<|x|<\delta\}\subset M_{\epsilon}.$
Recall $\epsilon<\delta^{2}\ll 1$. We use the identification map $x^{\prime}=\kappa(x)=x/\epsilon,\kappa^{*}(\alpha_{0})=\alpha^{\prime}$ to identify $g_{\epsilon}$ and $g_{\widehat{AH},\epsilon}$, and note that under this map $h_{\epsilon}$ and $h_{\epsilon^{\prime}}$ are identified up to $O(\epsilon^{3})$.
However in order to apply the analysis, we need to modify the space. The total space is glued from two parts: $\mathcal{W}_{0}$ corresponding to the Atiyah-Hitchin metrics in the center, and $\mathcal{W}_{1}$ corresponding to the multi-Taub-NUT singularities. Here $\mathcal{W}_{0}$ and $\mathcal{W}_{1}$ are essentially $\widehat{AH}_{\epsilon}$ and $M_{\epsilon}$ with the additional steps of compactification at infinity and singularity resolution at $\epsilon=0$. The construction of $\mathcal{W}_{1}$ and $\mathcal{W}_{0}$ are similar to [SS21] with a few modifications. See Figure 1 for an illustration.
Before we talk about the construction, we will first briefly recall the idea of real blow ups. For a general introduction the readers are referred to [Mel96] and for the specific construction for this paper we referred to [SS21, Section 3.2]. Given $X$ a manifold with corners, we call a submanifold $Y\subset X$ a $p$-submanifold if around any point $q\in Y$, there is a neighborhood ${\mathcal{U}}\subset X$ with a product decomposition of the form $(Y\cap{\mathcal{U}})\times[0,\epsilon)^{\ell-\ell^{\prime}}\times(-\epsilon,\epsilon)^{m-m^{\prime}}$. If $Y$ is a $p$-submanifold of $X$, we define the blowup $[X;Y]$ as the disjoint union of $X\setminus Y$ and $SN^{+}Y$, the inward-pointing spherical normal bundle of $Y$, with a smooth structure generated by the lifts of all smooth functions on $X$ and the spherical normal coordinates around $Y$. This blowup has a new front face equal to the set of all spherical normal vectors. Any vector field $V$ on $X$ lifts to a vector field on $[X;Y]$; this lift is nonsingular if and only if $V$ is tangent to $Y$. Blowup has the effect of making such tangent vector fields less degenerate.
Now we introduce the construction. The construction is first done on the double cover $\widetilde{\mathcal{W}}_{i},\ i=0,1$, and the final space $\mathcal{W}$ is the quotient by the involution $\iota$. First we construct $\widetilde{\mathcal{W}}_{0}$, which is defined as
@@
The construction of $\widetilde{\mathcal{W}}_{1}$ is the same as in [SS21, Section 4.1], obtained by pulling back the product $0,\delta^{2})_{\epsilon}\times M_{\epsilon}$ to the resolved base $\widetilde{B}$, where $\widetilde{B}$ is obtained through the blow up
(11) $\widetilde{B}=[\bar{\mathbb{R}}^{3}_{x}\times[0,\delta^{2})_{\epsilon};\{(0,0),\pm p_{i}\times\{0\},i=1,\dots,k-1\}]\cap\{|x|>\epsilon/\delta\}$
which introduces 2k12k-1 front faces. We then pull back the circle fibration $\phi:M_{\epsilon}\rightarrow\bar{\mathbb{R}}^{3}\setminus\{|x|\leq\epsilon/\delta\}$ to $\widetilde{B}$, such that when restricting to each $\epsilon>0$ it is the same as the original fibration $\phi$. On the other hand, when restricting to $\epsilon=0$ it lifts to a circle bundle over the blown up space $\bar{\mathbb{R}}^{3};\{0,\pm p_{i}\}]$.
We construct the total space by gluing $\widetilde{\mathcal{W}}_{1}$ into $\widetilde{\mathcal{W}}_{0}$, identifying the torus neighborhood using the lift of the map $\kappa$ described earlier, i.e. we identify the neighborhood
$\{(z^{\prime},\epsilon^{\prime}):0\leq\epsilon^{\prime}\leq\epsilon_{0},\delta^{-1}<|\psi(z^{\prime})|<\delta(\epsilon^{\prime})^{-1}\}\subset\widetilde{\mathcal{W}}_{0}$
@@
(12) $\widetilde{\mathcal{W}}:=\widetilde{\mathcal{W}}_{0}\coprod_{\kappa}\widetilde{\mathcal{W}}_{1},\ \mathcal{W}:=\widetilde{\mathcal{W}}/\iota.$
Here $\mathcal{W}$ is a manifold with corners, where the boundary faces include: $I_{\infty}$ which corresponds to the fibration over the boundary of the compactified $\mathbb{R}^{3}$ of $\mathcal{W}_{1}$; $X_{0}$ that is the lift of the fibration to the boundary face created by blowing up $(0,0)$; $X_{i},i=1,\dots k-1$, that are the lift of the fibration to the boundary faces created by blowing up $\{\pm p_{i},0\}$; $X_{ad}$ the remaining part of $\epsilon=0$. See Figure 1 for a picture of the total space.
Figure 1. An illustration of the resolved total space $\mathcal{W}$ obtained from gluing $\mathcal{W}_{0}$ and $\mathcal{W}_{1}$, with boundary faces and associated boundary defining functions on $\pi^{-1}(0)$.
There is a projection map $\pi:\mathcal{W}\rightarrow[0,\epsilon_{0})_{\epsilon}$, such that for each nonzero $\epsilon$ the fiber is an approximate ALF metric, while for $\epsilon=0$ the fiber is a manifold with corners. We also note that $\epsilon$ is the total boundary defining function, i.e. if we define $\sigma_{\nu}$ to be the boundary defining function of $X_{\nu},\nu=0,1,\dots,k-1$, and $\rho$ to be the boundary defining function of $X_{ad}$, then the product of all boundary defining functions on the front face is
@@ 2.4. Solving for the hyperKähler triples
By the gluing construction discussed above, we arrived at a family of approximate metrics, each living on the fiber $\pi^{-1}(\epsilon)$ of $\mathcal{W}$. We will follow the set up in  [SS21] to perturb these metrics and solve for the exact hyperKähler triples on each fiber. Recall that every hyperKähler metric on a 4-manifold gives a hyperKähler triple $\mathbf{\omega}=(\mathbf{\omega}_{1},\mathbf{\omega}_{2},\mathbf{\omega}_{3})$ such that each $\omega_{i}$ is a symplectic form and they satisfy the following equation [Don06]
@@
For an adiabatic Gibbons-Hawking ansatz in (1), the triple is given by the following $\mathbf{\omega}_{\epsilon}=(\mathbf{\omega}_{1},\mathbf{\omega}_{2},\mathbf{\omega}_{3})$ [SS21, (2.21)]
@@
which can be written as $\Omega_{\epsilon}+db$ where $b\in O(|\epsilon|^{3})$ and $\Omega_{\epsilon}$ is given by replacing $h_{\epsilon}$ by its leading term $1-\frac{\epsilon}{|x|}+\sum_{v=1}^{k-1}\frac{\epsilon}{|p_{v}|}$. Similarly for the branched cover $\widehat{AH}_{\epsilon^{\prime}}$ the triple is given by $\mathbf{\omega}_{AH,\epsilon^{\prime}}=\Omega_{\epsilon^{\prime}}+da$ where $a=O(e^{-|x^{\prime}|})$ and $\Omega_{\epsilon^{\prime}}$ is given by
@@
with
$h_{\epsilon}^{\prime}=1-\frac{1}{|x^{\prime}|}+\sum_{v=1}^{k-1}\frac{\epsilon}{|p_{v}|}.$
@@
Fix $\{0,p_{1},\dots,p_{k-1}\}$ a set of distinct $k$ points. There exists $\epsilon_{0}>0$, such that on $\pi^{-1}([0,\epsilon_{0}))\subset\mathcal{W}$, there exists smooth triple $\mathbf{\omega}_{Se}=(\mathbf{\omega}_{1},\mathbf{\omega}_{2},\mathbf{\omega}_{3})$ with $\mathbf{\omega}_{i}\in\Omega_{\phi}^{2}(\mathcal{W}/I)$ satisfying $Q(\mathbf{\omega})=0$. The corresponding fiber metric $g_{Se,\epsilon}$ satisfies
@@
    $g_{Se,0}|_{X_{0}}$ restricts to be the Atiyah–Hitchin metric, and $g_{Se,0}|_{X_{i}}$ restricts to be the Taub–NUT metric for $i=1,\dots,k-1$.
@@
where the two hyperKähler triples are defined in (15) and (16). Note that the harmonic function in $\mathbf{\omega}_{AH,\epsilon^{\prime}}$ and $\mathbf{\omega}_{\epsilon}$ agree up to order $\epsilon^{3}$, therefore the initial approximation solves the equation up to the following error
@@
The second step is to solve the equation iteratively to $\epsilon^{\infty}$ error. This step follows from the discussion in [SS21, Section 5], where they solve the error iteratively near the corner of $X_{\nu}\cap X_{ad}$ on the two faces. Recall that the linearized operator of $Q$ is given by 12 copies of scalar Laplacians by taking basis of self-dual 2-forms. When restricting to the solvability near $X_{\nu}$, note that the new component $AH$ is still a strongly ALF metric hence the analysis from [SS21] carries through. The solvability near $X_{ad}$ is the same as before. As a result we can iteratively construct a formal power series solution $\zeta=\mathbf{\omega}_{\chi}+dc$ such that
@@
The third step is to show there is a bounded inverse of the linearized operator on $\mathcal{W}$. This step again relies on that one can patch inverses on $X_{\nu}$ and $X_{ad}$. Since the only difference in the structure of $X_{\nu}$ from [SS21] is that $X_{0}$ is changed from $\overline{{\mathcal{M}}_{2}^{0}}$ by $\overline{AH}$, the inverse is constructed in the same way as in [SS21, Theorem C.3] for a strongly ALF space. On the other hand, on $X_{ad}$ the inverse is constructed explicitly using Fourier series and Green’s function where the only difference is the coefficient in $h_{\epsilon}$ which does not change the proof. As a result $\Delta_{\zeta}$ is an invertible operator on the domain constructed in [SS21, Theorem 6.6].
@@
As a corollary, such a family of metrics degenerates into the front face, and (under correct scaling and choice of base points) this gives Gromov-Hausdorff convergence to $AH$ and $TN$ respectively.
@@ 3. Non-holomorphic minimal spheres
In [MW93] Micallef and Wolfson showed that in the Atiyah–Hitchin space there is a special sphere that is stable minimal and not holomorphic with respect to any complex structures in the hyperKähler rotation. Using the gluing result above, we extend this result to $D_{k}$ spaces constructed above. As a perturbation result, the existence can be proved for $D_{k}$ spaces where the Taub-NUT singularities are sufficiently far from the origin.
@@
There exists $\epsilon_{0}$, such that for every $\epsilon\in(0,\epsilon_{0})$ there exists a stable minimal sphere in $g_{Se,\epsilon}$ that is non-holomorphic with respect to any of the compatible complex structures.
@@
We start with the approximate metric $g_{0,\epsilon}$ constructed via patching the $g_{\widehat{AH},\epsilon}$ and $g_{\epsilon}$. Since the minimal sphere in $g_{\widehat{AH}}$ is compact, so by shrinking $\epsilon_{0}$ if necessary, there exists a sphere in each fiber $\Sigma_{\epsilon}\subset(AH|_{|x^{\prime}|<\delta/\epsilon},g_{\widehat{AH},\epsilon})\subset(\pi^{-1}(\epsilon),g_{0,\epsilon})$ for every $\epsilon\in(0,\epsilon_{0})$. Note that this minimal sphere $\Sigma$ satisfies $\Sigma]\cdot[\Sigma]=-4$ which is not $2*\text{genus}(\Sigma)-2$, hence it cannot be holomorphic in a hyperKähler manifold via adjunct formula. Since the minimal sphere is stable in the Atiyah–Hitchin space [TW22], we can apply White’s implicit function theorem [Whi91] (which does not require the ambient manifold to be compact) to the actual hyperKähler metric $g_{SE,\epsilon}$ perturbed from $g_{0,\epsilon}$, we get a stable minimal non-holomorphic sphere in the $D_{k}$ space. ∎
@@
There is an open neighborhood in the moduli space of hyperKähler metrics on $K3$, such that each metric in this neighborhood contains a stable minimal sphere that is not holomorphic with respect to any compatible complex structures.
@@
The proof follows exactly the same as in [Fos19, Theorem 7.1] by considering any $D_{k}$ metric constructed above as a building block in the collapsing sequence. Since the approximate metric $g_{\epsilon}$ constructed using $D_{k}$ bubbles contains the special sphere, by the same perturbation argument in the previous theorem we obtain the existence of a non-holomorphic minimal sphere in a neighborhood. ∎
Remark.
We do not know whether the sphere we constructed is actually minimizing in those new ALF spaces and K3 surfaces. The model sphere in AH is area-minimizing, and in fact, calibrated [TW22]. It is an interesting question whether any of such properties will be preserved after the gluing construction.